OpenStudy (anonymous):

Factor square root of 296

OpenStudy (anonymous):

You want to look for numbers that go into 296 evenly that are squares. 16 is an example, which will not work here.

OpenStudy (anonymous):

2,4,8

OpenStudy (anonymous):

(296)^1/2=(4*74)^1/2=2(74)^1/2

OpenStudy (anonymous):

\[2 \sqrt{74} \]